Martinez Company’s relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows:

Average Cost Per Unit
Direct materials $ 6.10
Direct labor $ 3.60
Variable manufacturing overhead $ 1.40
Fixed manufacturing overhead $ 4.00
Fixed selling expense $ 3.10
Fixed administrative expense $ 2.10
Sales commissions $ 1.10
Variable administrative expense $ 0.55

1. For financial accounting purposes, what is the total amount of product costs incurred to make 10,000 units? (Do not round intermediate calculations.)

2. For financial accounting purposes, what is the total amount of period costs incurred to sell 10,000 units? (Do not round intermediate calculations.)

3. If 8,000 units are produced and sold, what is the variable cost per unit produced and sold? (Round your answer to 2 decimal places.)

4. If 12,500 units are produced and sold, what is the variable cost per unit produced and sold? (Round your answer to 2 decimal places.)

5. If 8,000 units are produced and sold, what is the total amount of variable costs related to the units produced and sold? (Do not round intermediate calculations.)

6. If 12,500 units are produced and sold, what is the total amount of variable costs related to the units produced and sold? (Do not round intermediate calculations.)

7. If 8,000 units are produced, what is the average fixed manufacturing cost per unit produced?

8. If 12,500 units are produced, what is the average fixed manufacturing cost per unit produced? (Round your answer to 2 decimal places.)

9. If 8,000 units are produced, what is the total amount of fixed manufacturing cost incurred to support this level of production?

10. If 12,500 units are produced, what is the total amount of fixed manufacturing cost incurred to support this level of production?

11. If 8,000 units are produced, what is the total amount of manufacturing overhead cost incurred to support this level of production? What is this total amount expressed on a per unit basis? (Round your "per unit" answer to 2 decimal places and other answers to the nearest whole dollar amount.)

12. If 12,500 units are produced, what is the total amount of manufacturing overhead cost incurred to support this level of production? What is this total amount expressed on a per unit basis? (Round your "per unit" answer to 2 decimal places and other answers to the nearest whole dollar amount.)

13. If the selling price is $22.10 per unit, what is the contribution margin per unit? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

14. If 12,000 units are produced, what are the total amounts of direct and indirect manufacturing costs incurred to support this level of production? (Do not round intermediate calculations.)

15. What incremental manufacturing cost will Martinez incur if it increases production from 10,000 to 10,001 units? (Round your answer to 2 decimal places.)
Answer #1

1) Total product cost on 10000 Units = (6.10+3.60+1.40+4) = 15.10*10000 = 151000

2) Total period cost on 10000 Units = (3.10+2.10+1.10+0.55)*10000 = 68500

3) Variable cost per unit sold = 6.10+3.60+1.40+1.1+0.55 = 12.75

4) Variable cost per unit sold = 12.75

Note: Please post whole question in the group of 4-4 sub parts as per HOMEWORKLIB RULES